# Towards Robust Spatio-Temporal Auto-Regressive Prediction: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

### v1

@@ Abstract
This study addresses the critical challenge of error accumulation in spatio-temporal auto-regressive predictions within scientific machine learning models by introducing innovative temporal integration schemes and adaptive multi-step rollout strategies. We present a comprehensive analysis of time integration methods, highlighting the adaptation of the two-step Adams-Bashforth scheme to enhance long-term prediction robustness in auto-regressive models. Additionally, we improve temporal prediction accuracy through a multi-step rollout strategy that incorporates multiple future time steps during training, supported by three newly proposed approaches that dynamically adjust the importance of each future step. By integrating the Adams-Bashforth scheme with adaptive multi-step strategies, our graph neural network-based auto-regressive model accurately predicts 350 future time steps, even under practical constraints such as limited training data and minimal model capacity—achieving an error of only 1.6% compared to the vanilla auto-regressive approach. Moreover, our framework demonstrates an 83% improvement in rollout performance over the standard noise injection method, a standard technique for enhancing long-term rollout performance. Its effectiveness is further validated in more challenging scenarios with truncated meshes, showcasing its adaptability and robustness in practical applications. This work introduces a versatile framework for robust long-term spatio-temporal auto-regressive predictions, effectively mitigating error accumulation across various model types and engineering discipline.
@@ 1 Introduction
In the era of scientific machine learning (SciML), auto-regressive (AR) models have emerged as powerful tools for spatio-temporal prediction across various engineering disciplines, particularly in physical domains like fluid dynamics [1, 2, 3, 4]. They predict future states by recursively using their own previous predictions as inputs for subsequent forecasts—specifically, each new prediction becomes part of the input sequence for the next prediction step, creating a chain of sequential forecasts based on historical data. This recursive approach facilitates real-time forecasting and dynamic decision-making across a wide range of engineering applications [5, 6, 7]. Unlike other emerging SciML approaches such as physics-informed neural networks or DeepONet, which incorporate the time coordinate directly into the input and therefore violate temporal causality [8, 9, 10], AR models predict future states based on past historical information. This sequential prediction aligns with the natural progression of physical processes, preserving causality by ensuring that each snapshot at a given time depends solely on preceding snapshots, without any influence from future information.
@@
However, AR models have inherent limitations, most notably error accumulation during long-term rollouts [2, 11, 12]. Since each prediction depends on the previous output, any inaccuracies introduced at one step can propagate and amplify in subsequent steps, leading to significant deviations from the true physical behavior over time. Recent approaches attempt to address this limitation by combining data-driven models with traditional numerical solvers, using the latter to recalibrate data-driven predictions when SciML model errors exceed certain thresholds [13, 14]. However, such hybrid approaches significantly compromise the primary advantage of ML-based surrogate models, their real-time prediction capability, as demonstrated by insufficient speedup factors: for example, approximately 1.9 times acceleration compared to pure computational fluid dynamics (CFD) simulations [14]. This highlights the critical need for improving the long-term prediction accuracy of purely data-driven AR models while preserving their computational efficiency, enabling real-time predictions without relying on expensive numerical solver re-calibrations.
@@
To enhance the robustness of AR-based spatio-temporal predictions over long-term rollouts, the most frequently used approach is noise injection, where noise is added to the input data during the training phase [12, 15, 16, 17]. This method adds random noise during training to help the model become robust against prediction errors that accumulate during the rollout phase. However, noise injection requires careful tuning of the noise scale, which is highly data-dependent, and its stochastic nature can lead to inconsistent and unstable training process.
@@
This study aims to present a novel framework for enhancing long-term AR predictions through the integration of numerical time-integration schemes and adaptive multi-step rollout techniques. Our approach specifically addresses the need for robust long-term prediction methods that maintain the real-time prediction capabilities of data-driven models while performing effectively under practical constraints such as limited datasets and computational resources. Extensive experiments, including comparisons with traditional noise injection methods and evaluations under challenging mesh conditions, have validated the model’s ability to achieve an optimal balance between prediction robustness and computational efficiency. Our key contributions can be summarized as follows:
@@
    Assessment under challenging but practical conditions: In real-world engineering scenarios, models often encounter challenging constraints such as limited availability of training data, restricted model capacity, and the need to handle varying mesh topologies arising from diverse shape configurations. To address these challenges, we aim to develop a GNN-based AR model that achieves accurate long-term predictions of future states (up to 350 rollouts) using only 50 past snapshots. Our approach emphasizes efficiency by utilizing a lightweight architecture with only 1,177 trainable parameters, specifically focusing on GNNs due to their inherent mesh-scalability and adaptability to diverse topologies.
@@
    Development of a comprehensive AR framework: We developed a comprehensive AR SciML framework that integrates traditional numerical schemes with advanced multi-step rollout approaches. We systematically evaluated multiple numerical schemes including forward Euler, first- and second-order central differences, and two-step Adams-Bashforth methods. Through extensive testing of these schemes coupled with different multi-step rollout strategies, we achieved significant improvements in prediction accuracy and robustness.
@@
    Enhancement through adaptive weighting strategies: To further improve the conventional multi-step rollout, we introduced three adaptive weighting strategies. Compared to the standard multi-step rollout with fixed weights, two of these strategies–—with learnable parameters–—demonstrated superior performance. Ultimately, combining the Adams-Bashforth method with an adaptive weighting approach that emphasizes only the first and last future components in the multi-step rollout process enabled accurate prediction of vortex shedding dynamics across all 350 rollouts.
@@
    Comparison with noise-injection methods: We validated our fusion of the Adams-Bashforth method with adaptive multi-step rollout against the conventional noise injection approach, commonly used to improve long-term rollout performance. Across various time schemes, our adaptive multi-step rollout consistently outperformed noise-injection methods.
@@
    Robustness evaluation on a more challenging mesh: To assess the robustness of our AR framework with GNN surrogate models, we conducted tests under even more challenging conditions using a truncated mesh that captures only the region behind the cylinder. The combination of the Adams-Bashforth method with adaptive multi-step rollout again significantly outperforms the conventional direct prediction AR approach with vanilla multi-step rollout.
@@
    General applicability of the proposed approach: Our versatile framework is designed to work beyond specific SciML models, such as GNNs, and specific applications, such as flow-field prediction. It provides a scalable and versatile solution applicable to any AR model across engineering disciplines dealing with spatio-temporal prediction tasks.
@@ 2.1.1 Motivation
In the realm of CFD, accurately predicting the temporal evolution of flow fields, particularly over extended time horizons, is crucial for understanding complex flow phenomena. While numerous advanced time-stepping schemes have been rigorously developed and studied in traditional CFD to enhance long-term prediction stability, their application in AI-driven flow prediction remains relatively unexplored. Most AI-based temporal flow prediction research has focused on direct prediction [1, 3, 4, 17] or simple forward Euler methods [2, 15, 16] for temporal integration, often struggling with error accumulation during extended AR rollouts. Consequently, the potential impact of other finite difference schemes on long-term prediction accuracy has largely been overlooked in this context. To address this gap, we propose pioneering the application of various numerical finite difference schemes to enhance long-term AR temporal prediction. Specifically, we integrate and systematically compare discretization techniques commonly used in finite difference methods to understand their behavior, strengths, and limitations within an AR framework, leveraging Graph U-Nets for fluid dynamics predictions over extended time horizons. This investigation represents the first comprehensive study to evaluate the stability and accuracy of diverse time integration schemes in the context of AR temporal prediction for fluid dynamics.
@@ 2.1.2 Investigated time integration schemes
In this study, we explore four different finite difference schemes for the time integration of Graph U-Net architecture and compare their performance with the conventional direct prediction approach. The objective is to assess how various finite difference numerical methods impact the accuracy and stability of the temporal evolution of the flow field. Note: The Runge-Kutta method, one of the most widely used time integration schemes in the CFD domain, is not included in this study due to its inherent incompatibility with the data-driven AR process. Further details are provided in A.
@@
$\mathbf{u}(t+\Delta t)=\text{AR}\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots}}\right)$ (1)
where $\mathbf{u}(t)$ represents the flow field at time $t$, $\Delta t$ is the time step size, $N$ is the number of past time steps used as input, and AR refers to the AR model with past $N$ snapshots as inputs. Here, it is important to note that $\Delta t=1$ can be assumed since $\Delta t$ represents the conceptual difference between successive snapshot datasets rather than an actual physical time interval [16]—therefore, $\Delta t=1$ is assumed for the simplicity throughout this study. Returning to the main point, this conventional approach does not explicitly leverage past snapshot datasets, even though they are used as inputs to black-box AR models, ultimately reducing both training and inference efficiency.
@@
Instead of predicting the next time step directly, forward Euler method [15, 18] makes the AR model predict the rate of change of the flow field:
$\mathbf{f}(t)=\text{AR}\big{(}\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)\big{)}$ (2)
where the rate of change $\mathbf{f}(t)$ is computed by:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}$ (3)
@@
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\mathbf{f}(t)$ (4)
During training, the label for $\mathbf{f}(t)$ is computed from the given data using the finite difference approximation shown in Eq. 3, and AR model is trained to output the corresponding $\mathbf{f}(t)$ as in Eq. 2. During inference, the AR model predicts $\mathbf{f}(t)$ based on past flow fields, and finally $\mathbf{u}(t+\Delta t)$ can be obtained as in Eq. 4.
2) First-order central difference
With this scheme, the AR model again predicts the $\mathbf{f}(t)$ as in Eq. 2 but the term $\mathbf{f}(t)$ is defined with the form of first-order central difference as [19]:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t-\Delta t)}{2\Delta t}$ (5)
@@
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t-\Delta t)+2\Delta t\cdot\mathbf{f}(t)$ (6)
@@
This method considers both forward and backward time steps for the calculation of $\mathbf{f}(t)$, potentially offering improved accuracy when flow properties vary smoothly over time. However, when the flow exhibits oscillatory behavior, this approach can lead to erratic predictions. Also, from Eq. 6, note that the next time step $\mathbf{u}(t+\Delta t)$ is highly dependent on $\mathbf{u}(t-\Delta t)$ without the intermediate time step $\mathbf{u}(t)$, which will be discussed later (Fig. 4 in Section 4.1).
3) Second-order central difference
Similar to the first-order central difference, this approach predicts $\mathbf{f}(t)$ but with a different definition using the second-order accurate central difference scheme [20]:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-2\cdot\mathbf{u}(t)+\mathbf{u}(t-\Delta t)}{\Delta t^{2}}$ (7)
and therefore the next time step can be calculated as:
$\mathbf{u}(t+\Delta t)=2\cdot\mathbf{u}(t)-\mathbf{u}(t-\Delta t)+\Delta t^{2}\cdot\mathbf{f}(t)$ (8)
Here, $\mathbf{f}(t)$ in the second-order scheme is defined symmetrically around the time step $\mathbf{u}(t)$ as in Eq. 7 using all three components: $\mathbf{u}(t-\Delta t)$, $\mathbf{u}(t)$, and $\mathbf{u}(t+\Delta t)$. Therefore, it can be expected to reduce the bias in the approximation, making it more reliable and often leading to more stable solutions than the first-order central difference, where there are only two components, $\mathbf{u}(t-\Delta t)$ and $\mathbf{u}(t+\Delta t)$.
4) Adams-Bashforth with forward Euler (Adams-Euler)
@@
Next, we explore the two-step Adams-Bashforth method, which requires derivative information for prediction. The general form of the Adams-Bashforth update is:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\left(\frac{3}{2}\mathbf{f}(t)-\frac{1}{2}\mathbf{f}(t-\Delta t)\right)$ (9)
where $\mathbf{f}(t)$ is the estimated derivative at time $t$. The method for computing the derivative, $\mathbf{f}(t)$, can be freely chosen by the user. In this study, we focus primarily on the forward Euler method for approximating the derivative $\mathbf{f}(t)$—the Adams-Bashforth with forward Euler method will be referred to as Adams-Euler for the rest of the paper. With this regard, AR model aims to predict the corresponding $\mathbf{f}(t)$:
@@
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}=\text{AR}\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots from $t$}}\right)$ (10)
Similarly, the previous derivative $\mathbf{f}(t-\Delta t)$ is computed as:
@@
$\mathbf{f}(t-\Delta t)=\frac{\mathbf{u}(t)-\mathbf{u}(t-\Delta t)}{\Delta t}=\text{AR}\left(\underbrace{\mathbf{u}(t-\Delta t),\mathbf{u}(t-2\Delta t),\dots,\mathbf{u}(t-N\Delta t)}_{\text{past $N$ snapshots from ($t-\Delta t$)}}\right)$ (11)
Since both $\mathbf{f}(t)$ and $\mathbf{f}(t-\Delta t)$ can be obtained through Eq. 10 and Eq. 11, the next snapshot can be computed as in Eq. 9. A key feature of the two-step Adams-Bashforth method is its reliance on derivative information from previous time steps, requiring the storage of past outputs ($\mathbf{f}(t-\Delta t)$) to compute future predictions. Due to its property, three successive snapshots, $\mathbf{f}(t-\Delta t)$, $\mathbf{f}(t)$, and $\mathbf{f}(t+\Delta t)$, can be explicitly considered for the prediction of the next snapshot. This approach is particularly well-suited for problems where incorporating derivative approximations from the past is essential for accurately capturing the system’s temporal evolution.

### v2

# Long-Term Auto-Regressive Prediction using Lightweight AI Models: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

@@ Abstract
This study addresses the critical challenge of error accumulation in spatio-temporal auto-regressive predictions within scientific machine learning models by introducing innovative temporal integration schemes and adaptive multi-step rollout strategies. We present a comprehensive analysis of time integration methods, highlighting the adaptation of the two-step Adams-Bashforth scheme to enhance long-term prediction robustness in auto-regressive models. Additionally, we improve temporal prediction accuracy through a multi-step rollout strategy that incorporates multiple future time steps during training, supported by three newly proposed approaches that dynamically adjust the importance of each future step. Despite using an extremely lightweight graph neural network with just 1,177 trainable parameters and training on only 50 snapshots, our framework accurately predicts 350 future time steps—a 7:1 prediction-to-training ratio—achieving an error of only 1.6% compared to the vanilla auto-regressive approach. Moreover, our framework demonstrates an 83% improvement in rollout performance over the standard noise injection method, a standard technique for enhancing long-term rollout performance. Its effectiveness is further validated in more challenging scenarios with truncated meshes, showcasing its adaptability and robustness in practical applications. This work introduces a versatile framework for robust long-term spatio-temporal auto-regressive predictions that shows potential for mitigating error accumulation across various model types and engineering disciplines.
@@ 1 Introduction
In the era of scientific machine learning (SciML), auto-regressive (AR) models have emerged as powerful tools for spatio-temporal prediction across various engineering disciplines, particularly in physical domains like fluid dynamics [1, 2, 3, 4, 5, 6]. They predict future states by recursively using their own previous predictions as inputs for subsequent forecasts—specifically, each new prediction becomes part of the input sequence for the next prediction step, creating a chain of sequential forecasts based on historical data. This recursive approach facilitates real-time forecasting and dynamic decision-making across a wide range of engineering applications [7, 8, 9, 10, 11, 12]. Unlike other emerging SciML approaches such as physics-informed neural networks or DeepONet, which incorporate the time coordinate directly into the input and therefore violate temporal causality [13, 14, 15], AR models predict future states based on past historical information. This sequential prediction aligns with the natural progression of physical processes, preserving causality by ensuring that each snapshot at a given time depends solely on preceding snapshots, without any influence from future information.
@@
However, AR models have inherent limitations, most notably error accumulation during long-term rollouts [2, 16, 17, 18, 19]. Since each prediction depends on the previous output, any inaccuracies introduced at one step can propagate and amplify in subsequent steps, leading to significant deviations from the true physical behavior over time. Recent approaches attempt to address this limitation by combining data-driven models with traditional numerical solvers, using the latter to recalibrate data-driven predictions when SciML model errors exceed certain thresholds [20, 21]. However, such hybrid approaches significantly compromise the primary advantage of ML-based surrogate models, their real-time prediction capability, as demonstrated by insufficient speedup factors: for example, approximately 1.9 times acceleration compared to pure computational fluid dynamics (CFD) simulations [21]. This highlights the critical need for improving the long-term prediction accuracy of purely data-driven AR models while preserving their computational efficiency, enabling real-time predictions without relying on expensive numerical solver re-calibrations.
@@
To enhance the robustness of AR-based spatio-temporal predictions over long-term rollouts, the most frequently used approach is noise injection, where noise is added to the input data during the training phase [17, 22, 23, 24]. This method deliberately perturbs the training inputs with random noise to simulate the prediction errors that naturally occur during rollout. By exposing the model to slightly corrupted inputs during training, it learns to handle imperfect data and becomes more robust when its own imperfect predictions are fed back as inputs during inference. However, noise injection requires careful tuning of the noise scale, which is highly data-dependent, and its stochastic nature can lead to inconsistent and unstable training process.
@@
This study presents a novel framework for enhancing long-term auto-regressive (AR) predictions in resource-constrained environments through the integration of numerical time-integration schemes and adaptive multi-step rollout techniques. Our approach specifically addresses a critical challenge in scientific machine learning: achieving robust long-term predictive performance with limited training data and model capacity while maintaining real-time inference capabilities. Extensive experiments, including comparisons with traditional noise injection methods and evaluations under challenging mesh conditions, demonstrate our framework’s ability to achieve exceptional long-term stability and accuracy despite these constraints. Our key contributions can be summarized as follows:
@@
    Assessment under challenging but practical conditions: In real-world engineering scenarios, models often encounter challenging constraints such as limited availability of training data, restricted model capacity, and the need to handle varying mesh topologies arising from diverse shape configurations. To address these challenges, we develop a GNN-based AR model that achieves accurate long-term predictions of future states (up to 350 rollouts) using only 50 past snapshots, with a lightweight architecture containing only 1,177 trainable parameters. This demonstrates the feasibility of robust long-term prediction even under severe resource constraints that are common in practical applications.
@@
    Comprehensive evaluation of time integration schemes for AR prediction: For the first time in the context of AR prediction within SciML models, we comprehensively explore various time integration schemes, including direct prediction, forward Euler, second-order central difference, and two-step Adams-Bashforth methods. Through systematic evaluation, we demonstrate that the Adams-Bashforth scheme consistently outperforms other approaches, achieving a 7% improvement in prediction accuracy compared to the commonly used forward Euler method. While other schemes exhibit instability with increased multi-step rollout, Adams-Bashforth maintains robust performance across varying prediction horizons, even when the direct prediction approach fails to converge.
@@
    Proposal of adaptive multi-step rollout strategies: We introduce a novel concept of adaptive multi-step rollout to address limitations in conventional multi-step training approaches. By developing three adaptive weighting strategies that dynamically adjust the importance of different future time steps, we significantly enhance prediction stability. Our best approach, which emphasizes only the first and last future components, delivers an 89% improvement over conventional fixed-weight multi-step rollout and outperforms other adaptive strategies by up to 80%, while maintaining similar computational costs. This demonstrates the effectiveness of selectively emphasizing critical time steps during training.
@@
    Comparison with noise-injection methods: Our integration of the Adams-Bashforth method with adaptive multi-step rollout achieves an 83% improvement in prediction accuracy compared to conventional noise injection techniques—the current standard approach for enhancing long-term rollout stability. Our analysis reveals that combining multi-step rollout with noise injection degrades performance due to a previously unidentified negative interaction between these error-mitigation strategies.
@@
    Robustness evaluation on a more challenging mesh: Tests using a truncated mesh—where only a subset of the complete mesh containing all flow quantities is used for training—show that our method achieves 58% and 27% improvement over direct prediction and forward Euler approaches respectively. This demonstrates our framework’s applicability in scenarios where computational memory constraints necessitate focusing on specific regions of interest within complex flow fields.
@@
    General applicability of the proposed approach: We intentionally design our framework to be model-agnostic and application-independent, focusing on fundamental AR prediction mechanics rather than domain-specific features. This ensures our techniques can be seamlessly integrated with any AR model across diverse engineering disciplines without requiring specialized adaptations.
@@ 2.1.1 Motivation
While advanced time-stepping schemes have been extensively studied in traditional CFD, their application in AI-driven flow prediction remains limited. Most AI-based temporal flow prediction research has focused only on direct prediction [1, 3, 4, 24] or simple forward Euler methods [2, 22, 23], leaving the potential impact of other finite difference schemes largely unexplored.
To address this gap, we investigate various numerical finite difference schemes for enhancing long-term AR prediction accuracy. Unlike traditional numerical analysis, where schemes are evaluated based on formal order-of-accuracy, AR models implicitly learn temporal dynamics from data. In our framework, each finite difference method computes the training target for temporal derivatives with inherently different structural biases—from simple first-order approximations in forward Euler to symmetric formulations in central difference schemes to the multi-step approach of Adams-Bashforth which incorporates richer temporal history.
This investigation represents the first comprehensive study to evaluate how different derivative approximation methods affect AR model’s ability to capture spatiotemporal dynamics in fluid flow prediction. The observed benefits with particular integration methods arise from favorable interplay between numerical approximation and the model’s data-driven learning process.
@@ 2.1.2 Investigated time integration schemes
We explore three finite difference schemes (forward Euler, second-order central difference, and Adams-Bashforth) and compare them with conventional direct prediction. It’s important to note that the formal order-of-accuracy concept doesn’t directly translate to AR models, as they learn temporal dependencies implicitly rather than following fixed error bounds. Figure 1 visualizes model outputs across all schemes. The Runge-Kutta method, though widely used in CFD, is excluded due to its incompatibility with data-driven AR processes (details in A).
Figure 1: Overview of time integration schemes investigated in this study. While both forward Euler and Adams-Bashforth yield identical model predictions during simulation, they differ in how they compute the subsequent timestep $u(t+\Delta t)$ during post-processing: forward Euler uses a single-step explicit update (Eq. 4), whereas Adams-Bashforth employs a two-step approach incorporating previous timesteps (Eq. 7).
@@
$\mathbf{u}(t+\Delta t)=AR\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots}}\right)$ (1)
where $\mathbf{u}(t)$ represents the flow field at time $t$, $\Delta t$ is the time step size, $N$ is the number of past time steps used as input, and $AR$ refers to the AR model with past $N$ snapshots as inputs. Here, it is important to note that $\Delta t=1$ can be assumed since $\Delta t$ represents the conceptual difference between successive snapshot datasets rather than an actual physical time interval [23]—therefore, $\Delta t=1$ is assumed for the simplicity throughout this study. Returning to the main point, this conventional approach does not explicitly leverage past snapshot datasets when predicting $\mathbf{u}(t+\Delta t)$, even though they are used as inputs to black-box AR models, ultimately reducing both training and inference efficiency.
@@
Instead of predicting the next time step directly, forward Euler method [22, 25, 26, 27] makes the AR model predict the rate of change of the flow field, $\frac{\delta\mathbf{u}}{\delta t}$:
$\frac{\delta\mathbf{u}}{\delta t}=AR\big{(}\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)\big{)}$ (2)
where the rate of change is computed by:
$\frac{\delta\mathbf{u}}{\delta t}=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}$ (3)
@@
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\frac{\delta\mathbf{u}}{\delta t}$ (4)
Before training, the label for $\frac{\delta\mathbf{u}}{\delta t}$ is computed from the given data using the forward Euler method shown in Eq. 3, and AR model is trained to output the corresponding $\frac{\delta\mathbf{u}}{\delta t}$ as in Eq. 2. During inference, the AR model predicts $\frac{\delta\mathbf{u}}{\delta t}$ based on past flow fields, and finally $\mathbf{u}(t+\Delta t)$ can be obtained as in Eq. 4.
2) Second-order central difference
With this scheme, the AR model again predicts the $\frac{\delta\mathbf{u}}{\delta t}$ as in Eq. 2 but the term $\frac{\delta\mathbf{u}}{\delta t}$ is defined with the form of second-order central difference [28] as:
$\frac{\delta\mathbf{u}}{\delta t}=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t-\Delta t)}{2\Delta t}$ (5)
@@
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t-\Delta t)+2\Delta t\cdot\frac{\delta\mathbf{u}}{\delta t}$ (6)
@@
This method considers both forward and backward time steps for the calculation of $\frac{\delta\mathbf{u}}{\delta t}$, potentially offering improved accuracy when flow properties vary smoothly over time. However, when the flow exhibits oscillatory behavior, this approach can lead to erratic predictions. Also, from Eq. 6, note that the next time step $\mathbf{u}(t+\Delta t)$ is highly dependent on $\mathbf{u}(t-\Delta t)$ without the intermediate time step $\mathbf{u}(t)$, which will be discussed later (Fig. 5 in Section 4.1).
3) Adams-Bashforth with forward Euler (Adams-Euler)
@@
Next, we explore the two-step Adams-Bashforth method, which again requires derivative information for prediction. The general form of the Adams-Bashforth update is:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\left(\frac{3}{2}\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}-\frac{1}{2}\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}\right)$ (7)
where $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}$ represents the rate of change of $\mathbf{u}$ at the current timestep and $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}$ denotes the rate of change at the previous timestep.
A critical aspect of implementing the Adams-Bashforth method in an AR framework is the choice of how to approximate the derivatives $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}$ and $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}$. We implement the forward Euler method for time integration (Eq. 8 and Eq. 9), a common approach in AI-based temporal prediction [2, 22, 23]. Since both methods employ the forward Euler scheme to compute the derivatives, the forward Euler and Adams-Bashforth approaches yield identical model outputs: see Eq. 1 and Eq. 8. For brevity, we refer to the combined Adams-Bashforth with forward Euler method as “Adams-Euler” throughout this paper.
@@
$\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}=AR\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots from $t$}}\right)$ (8)
@@
$\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}=\frac{\mathbf{u}(t)-\mathbf{u}(t-\Delta t)}{\Delta t}=AR\left(\underbrace{\mathbf{u}(t-\Delta t),\mathbf{u}(t-2\Delta t),\dots,\mathbf{u}(t-N\Delta t)}_{\text{past $N$ snapshots from ($t-\Delta t$)}}\right)$ (9)
The Adams-Euler method thus comprises two main components:
  1. 1.
    Forward Euler derivative approximation: We compute the derivatives $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}$ and $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}$ using the forward Euler method (Eqs. 8 and 9). This approach is consistent with previous AR studies and provides a straightforward and effective means to compute derivatives, which will be used as training labels [22, 23].
  2. 2.
    Two-step Adams-Bashforth integration: The update in Eq. 7 utilizes both $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}$ and $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}$ to predict $u(t+\Delta t)$, thereby integrating historical derivative information. This key distinction differentiates Adams-Euler from forward Euler, even though both methods produce identical model outputs, $\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}$.